How many mililiters of a 10% HCl solution and a 30% HCl solution should be mixed to make 200 mililiters of a 15% HCl solution? Hurry! I have to go to class soon! Thanks for the help!
Answer by rothauserc(4718)

x + y = 200
.10x + .30y = .15*200 = 30
solve first equation for x
x = 200 -y
now substitute for x in second equation
.10*(200 - y) + .30y = 30
20 -.10y + .30y = 30
.20y = 10
y = 50
x = 150
check answers in second equation
.10*150 + .30*50 = 30
15 + 15 = 30
30 = 30
answers check
therefore we need 150 ml of 10% HCI solution and 50 ml of 30% HCI solution